Morrow Corporation had only one job in process during May—Job X32Z—and had no finished goods inventory on May 1. Job X32Z was started in April and finished during May. Data concerning that job appear below: Job X32Z Beginning balance $ 6,100 Charged to the job during May Direct materials $ 10,900 Direct labor $ 3,100 Manufacturing overhead applied $ 5,100 Units completed 210 Units in process at the end of May 0 Units sold during May 95 In May, overhead was overapplied by $410. The company adjusts its cost of goods sold every month for the amount of the overhead that was underapplied or overapplied. Required: 1. Using the direct method, what is the cost of goods sold for May? 2. What is the total value of the finished goods inventory at the end of May? 3. What is the total value of the work in process inventory at the end of May?
Answer #1
STATEMENT OF COST OF GOODS MANUFACTURED
Beginning work in process 6100
Cost incurred during may
Direct material 10900
Direct labor 3100
Manufacturing overhead applied 5100
Total cost incurred during may 19100
Total work in process 25200
less:Ending work in process (0)
cost of goods manufactured 25200
units manufactured 210
Cost per unit [25200/210 ] $ 120 per unit

1)

cost of goods sold = Units sold * cost per unit

                      = 95*120

                       = $ 11400

Adjusted cost of goods sold =Unadjusted cost of goods sold + underapplied/(overapplied) overhead

                         = 11400+ (-410)

                            = 11400-410

                             = $ 10990

2)Finished goods inventory at end = Beginning +units completed- units sold

                      = 0+ 210 - 95

                       = 115 units

cost of finished goods inventory at end = cost per unit *Finished goods inventory at end

                                      = 120 * 115

                                        = $ 13800

3) the total value of the work in process inventory at the end of May = 0

since there is no work in process inventory at end of may